Required information [The following information applies to the questions displayed below.) Finlon Upholstery, Inc. uses a job5-a. Compute the amount of under- or overapplied overhead at year-end.5-b. Prepare the necessary journal entry to record its disposition. (If no entry is required for a transaction/event, select

1 Record under- or over applied overhead.
Answer #1

1) Overhead rate = 5460000/4200000 = 130% of labor cost

Applied overhead = 4350000*1.3 = 5655000

Actual overhead = 64000+2860000+1740000+58000+828000 = 5550000

Over applied overhead = 5655000-5550000 = 105000

Journal entry

Transaction General Journal Debit Credit
1 Manufacturing overhead 105000
Cost of goods sold 105000